SHOW ALL OF YOUR WORK

6. A firm has cash of 100,000, accounts receivable of 75,000, prepaid expenses of 12,500, accounts payable of 50,000, other current liabilities of 35,000, common stock of 375,000 and long term liabilities of 65,000. The firm also produced a profit of 20,000 during the last calendar year. What is the firms working capital?   
7. Which company has the most risk from an investor's standpoint? Firm A has a PE of 84 times and Firm B has a PE of 16 times. Assume both firms operate in the same industry. Firm A has fewer shares outstanding than Firm B.

A.

Firm A because it has the higher PE ratio.

B.

Firm B because it has a lower PE ratio.

C.

Firm A because it has fewer shares outstanding.

D.

Firm B because it has more shares outstanding.

8. A firm has an ACP of 45 days and its annual sales are $5.3 million. What is its account receivable balance?

9. An investor wanting large returns will be interested in companies that have:

A.

high ROAs.

B.

high ROEs.

C.

high current ratios.

D.

high times interest earned.

10. Which type of ratio measures a firm's ability to pay off short-term obligations without relying on inventory sales?
Answer #1

Question 6.

Working capital = Current assets – Current liabilities

Current assets = cash + accounts receivables + prepaid expenses

Current assets = 100000+75000+12500 = $ 187,500

Current liabilities = Accounts payable + other current liabilities

Current liabilities = 50000+35000 = 85000

Therefore, Working capital = 187500- 85000 = $ 102,500

Question 7

Firm A P/E ratio = 84 times whereas Firm B P/E ratio = 16 times

The firm with lower PE ratio is more stable than a firm with higher PE ratio

So as per above discussion Firm A is risky then Firm B due to higher PE ratio.

Question 8

ACP = 45 days

Annual sales = $ 5.3 million

ACP = 365 / receivable turnover ratio

So, Receivable turnover ratio = 365/ 45 = 8.11

Receivable turnover ratio = Sales revenue / Avg Accounts receivable

So, Accounts receivable = 5.3 million / 8.11 = $ 653,424.66

Question 9

Investor who requires higher returns from the stocks would invest in a company which has more ROEs. As if the stock is having more ROEs it would have the capacity to pay more returns.

Question 10

Quick ratio or Acid test ratio is the ratio that measures the firm’s ability to pay short obligations without relying on the inventory turnover.